OpenStudy (anonymous):

solve for b. (5b-2)^2 + 2(5b-2) - 24 =0

OpenStudy (anonymous):

let 5b-2 = y equation becomes: y^ 2+2y-24 =0 =>y =-6 or +4 =>5b-2 =-6 or 5b-2 =4 =>b=-4/5 or b=6/5